40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7350     Entry time: Thu Sep 6 16:46:44 2012
Author: Jenne 
Type: Update 
Category: SUS 
Subject: BS aligned, target removed 

Q and I aligned the BS such that we were hitting the center of ETMX. The ETMX cage does not have OSEM setscrew holes on the front, so it is not possible to put the targets that Steve made on this optic.  So, I put the freestanding ruler in front of the optic, with the edge of the ruler at the center (as viewed from above) of the optic.  Then Eric steered the BS until we were hitting the 5.5" mark, and roughly half of the beam was obscured by the ruler.

We then aligned ITMX such that the prompt reflection was colinear with the incoming beam. 

I checked the 2 spots through the BS, heading to the AS port.  (2 spots since MICH hasn't been locked / finely aligned yet).  They were being clipped on the 2nd output PZT.  I adjusted the knobs of the first output PZT to center the spots on the 2nd PZT.  Note that the output PZTs' power is still off, and has been off for some unknown length of time.  I had found them off when prepping for the vent a week or two ago.  So the current alignment depends on them staying off.  We don't really need them on until we're ready to employ our OMC.

The beams now look nicely unclipped on the AS camera, and we're aligning MICH.

ELOG V3.1.3-